ACIS Thermal/Vacuum Test Page

Latest update: Fri Apr 4 14:31:41 1997 EST

ACIS Home Page Status Plan History Notes Anomalies Phone Numbers Shifts
Live Engineering Data Thermal/Vacuum Procedure Long-Form Procedure Short-Form Procedure

Current Status:

  • Instrument in stasis>

    Short Term Plan:


    Noise Test Plan


    Test Event History

    Day Event Results
    2/11, Tue. Ship to Lincoln

    Prepare SIM-SIM Chamber leak checked
    2/12, Wed. Install SIM-SIM Done
    2/13, Thu. Troubleshoot the EGSE
    2/14, Fri. Troubleshoot for noise;
    Acceptable noise environment, but it will not be convenient for the science tests.

    Pump down Started at 1930
    2/15, Sat. First Long Form OK

    Open DA Valve Initial internal pressure was 14 torr; unit had last been pumped the previous Monday.
    2/16, Sun. Cold Soak #1 Short Form OK, but noisy
    2/17, Mon. Hot Soak #1 Short Form OK
    2/18, Tue. Cold Soak #2 Short Form OK; Bake Out Mode
    2/19, Wed. Hot Soak #2 Short Form OK; Bake Out Mode

    DH Door Opened Run LF Valve and Door Cycles

    Flight Software 1.1 Loaded Took about 4 hours with LRCTU special program; burned both side simultaneously.
    2/20, Thu. Cold Soak #3 Short Form OK; Doing Science

    Hot Soak #3 Short Form OK; Doing Science
    2/21, Fri. Cold Soak #4 Short Form OK; Doing Science

    Hot Soak #4 Long Form OK
    2/22, Sat. Cold Soak #5 Short Form OK; Doing Science

    Hot Soak #5 Short Form OK; Doing Science
    2/23, Sun. Cold Soak #6 Short Form OK; Doing Science

    Hot Soak #6 Short Form OK; Doing Science
    2/24, Mon. Cold Soak #7 Short Form OK; Doing Science

    Hot Soak #7 Short Form OK; End of Science for present
    2/25, Tue. Cold Survival; Cold Soak #8 Long Form OK
    2/26, Wed. DH Door and Valves Closed Normal Operation at 25C DH temperature
    2/27, Thu. 1238 Pre-Bake
    2/28, Fri. 1238 OWS Exposure Started exposure at 5am.
    3/1, Sat. Repressurize; Remove OWS OWS ooks bright and shiny

    Post T/V Long Form Functional OK

    Pump Down Started at 3:30pm
    3/2, Sun. Cooling to normal operating mode
    3/3, Mon. Noise/science testing
    3/4, Tue. Noise/science testing
    3/5, Wed. Science testing
    3/6, Thur. Science tests; close door Never did see internal LED
    3/7, Fri. Door and valves closed at 8am
    3/8, Sat. Repressurize; test with "fixed" Flight PSMC No joy.




    Test Notes

    • The DEA telemetry continues to be noisy.
    • The PSMC telemetry for the DEA Current is very noisy.

      Anomalies, Major and Minor:

      • At the 40 electron level, we are seeing noise on the CCD data. It has a 2KHz ringing characteristic we have seen before, and seems to be synchronous with the pixel clock! It is weakly dependant upon the number of DEA boards on and strongly dependant upon the number of DPA boards on. We'll run the Short Forms with higher clip levels,
      • The low conductance valves (both were opened) took 1 hour to reduce internal pressure from 14 torr to 1.5 torr; the Flight unit will take a fair amount longer.
      • We tripped the 28VDC Power Supply when we went into Bake Out Mode; all the FEPs were on. This is a message for the spacecraft folks, originally delivered a year ago. Specified thermal power levels don't translate into what someone might command the instrument to do. Turning on FP and DA Bake Out heaters, at 22VDC, adds 5.0A to our current drain. (And normal ops at 22VDC is only 7A.)
      • We're only getting down to -123C on the Cold Radiator -- and this in the cold case; it bespeaks of a heat leak. . . one which we will have to address.
      • During Cold Long Form, First Short Form, by accident commanded on both DEA power supplies and on DPA power supply. It should not have been possible, per LMA description, but both DEA supplies came up and delivered regulated voltages to both cards 11 and 12 in the DEA. It is good that we built safeties into !! and 12 against this occurance! (If DEA-A is on, -B is locked out, but not vice versa.)
      • During Cold Long Form, the B Side Door Drive kicked out on overtemp during the Door Close. Door did not fully close according to the serial data from the PSMC. Worked oK back at 25C.
      • The CCDs have in their noise spectrum a large 2KHz component. That component is synchronized to the pixel clock when the PSMC is warm (chassis and active DEA card above 0C), and unsynchronized when the PSMC is cooler. It is assumed that the PSMC is losing lock with the supplied 100KHz from the DPA in the cooler case (and this, given the magnitude of the noise, is in the "gotta fix" category).
      • If main S/C power is brought up very slowly, the PSMC DH heater circuits come up Enabled (maybe On, too, but we were to hot to tell).
      • Internal LED is a mystery. It worked with a test box, but neither the CCDs nor the eyeball could see the light.
      • There is a substantive light leak between the optical blocking filters; Pete will generate an ECO.

        What We've Learned about Noise:

        • There is a 2KHz problem and a 10-40KHz problem.
        • The 2KHz goes out of frame sync when either a) the PSMC DEA board is less than 0C, or b) at +5C the input bus voltage exceeds 34VDC (athough the noise spectra are not quite the same in the two cases).
        • The 10-40KHz noise is dependant solely upon FEP activity.
        • DEA board parallel clock transfers have nothing to do with either noise source.
        • Stopping the FEP from data processing has no affect on the 2KHz.
        • The EU PSMC does not exhibit either of the above discrete noise problems when run with EU DPA/DEA electronics; it also does not have CS06 problems.
        • Changing the jumper on the DPA feedback (within the EU PSMC) significantly reduces RMS noise as well as CS01 problems.

          Phone Numbers

          (from MIT, Lincoln Lab is 181 + 4 digits)

          (from Lincoln Lab, MIT is 806 + 5 digits)

          Name Function Work Home Notes
          Ellen Sen Ultimate TC 253-8342 617-665-3568
          Demitrios Athens Software GSE 258-5285 617-566-7668
          Mark Bautz Proj. Science 253-0023 617-863-2024
          Mike Doucette Electrical GSE 253-3571 617-438-7690
          Bob Goeke Proj. Eng. 253-1910 617-729-8484
          Brian Klatt Quality 253-7555 508-785-1511
          Bill Mayer Proj. Mgr. 253-7552 508-692-3318
          Matt Smith Contamination 252-1736 508-435-9970
          Ed Boughan TC 253-6146 508-470-1043
          Mary Briggs TC 258-7896 508-372-0358
          Sean Callaghan TC
          617-969-6178
          Peter Ford TC 253-7277 617-661-0868
          Gordon Gong TC 253-8135 617-969-3768
          Dorothy Gordon TC 253-0797 617-232-7327
          Hans Govaert TC 253-1977 508-535-1467
          John Hughes TC 253-8009 508-497-8344
          Fred Kasparian TC 253-1065 617-327-6330
          Jim O'Connor TC 253-6139 617-335-2915
          Joanne Vining TC 258-8090 617-396-9737
          Engineering GSE TC sits here 981-4604

          Science GSE
          981-2010

          Cleanroom/TV
          981-5138

          Ron Efromson LL Coordinator 981-4513

          Mike Wirzbicki LL Contamination 981-4326 981-5348 (beep)617-430-1886
          Jon Howell LL Vacuum 981-4990
          (beep)617-430-5684
          Angelo Calao LL Manager 981-7717


          Engineering GSE Shifts

          Day 12mid-8am 8am-4pm 4pm-12mid
          3/8, Saturday Callaghan Vining Mayer
          3/9, Sunday Callaghan O'Connor Hughes
          3/10, Monday Callaghan Briggs Klatt
          3/11. Tusday Callaghan Vining Gpvaert
          3/12, Wednesday Callaghan


          Page maintainer:
          Bob Goeke.

          There have been 170 accesses to this page since 9/17/2022.